True or False:

( ) As defined by Goldberger (1972), a structural model represents causal relationships.

( ) Considering the model y=a+bx+u E[u]=0 is a sufficient condition for estimating the parameter b consistently.

( ) Considering the model y=a+bx+u if C[x,u]=0 we can say that x is "exogenous" in the model.

( ) Considering the model y=a+bx+u the normal distribution of u is not necessary for the validity of the t-test and of the F-test, provided that the sample is sufficiently large.
Answer #1

TRUE:As defined by Goldberger (1972), a structural model represents causal relationships.

FALSE:Considering the model y=a+bx+u E[u]=0 is a sufficient condition for estimating the parameter b consistently.

TRUE: Considering the model y=a+bx+u if C[x,u]=0 we can say that x is "exogenous" in the model.

FASLE:

Considering the model y=a+bx+u the normal distribution of u is not necessary for the validity of the t-test and of the F-test, provided that the sample is sufficiently large.

THANKS

revert back for doubt

please upvote